OpenStudy (anonymous):

Use estimation to determine which of the following quotients you expect to be less than 1. thirteen-fifteenths divided by two-fifteenths seven-eighths divided by one-eighth eight-elevenths divided by two-elevenths three-sevenths divided by six−sevenths

Directrix (directrix):

I would go with the option that has a numerator that is smaller than its denominator. Do you see an option that looks like that?

Directrix (directrix):

@messiahcollins Look at this option: three-sevenths divided by six-sevenths. 3/7 is smaller than 6/7 so the quotient should be less than one. Let's see. 3/7 divided by 6/7 = 3/7 * 7/6 = 3/6 = 1/2. 1/2 < 1